# Gravity as a deformed topological gauge theory

### v6

@@ Abstract
Two examples are studied, Lorentzian geometry (including dS and AdS spacetimes) and Lorentz×$\times$×Weyl geometry. In the latter case, we prove the equations of motion exhibit a secondary source for curvature in addition to the usual energy-momentum tensor. This additional source is expressed in terms of the spin density of matter, torsion and their variations.
@@
Keywords: Topological gravity, Homogeneous spaces, Cartan geometry, Characteristic classes
1 Introduction
@@
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (3)
@@
Definition 1.1.
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a 111-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [2, p128]:
@@
$(b)~(R_{h})^{*}\varpi=Ad(h^{-1})\varpi~,\forall h\in H,\text{ where }R\text{ is the right action of }H\text{ on }\mathscr{P}$
$\quad\ \,\text{and }Ad\text{ the adjoint representation of }H\subset G\text{ on the Lie algebra }\mathfrak{g}.$
@@
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$. This action is $Ad(H)$-invariant, i.e. Lorentz invariant.
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$ originating from the $\mathfrak{so}(m)$-valued connection 111-form $\omega_{\mathfrak{so}}$. This is an $SO(m)$-invariant polynomial.
For a 444-dimensional manifold one has:
@@
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$, and $-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$ acts as the deformation of the $\bar{\Omega}_{\mathfrak{so}}=R$ curvature.
@@
If $\mathfrak{g}/\mathfrak{h}=\mathfrak{m}$ obeys the commutation relation $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ (symmetric Lie algebra), then $G/H$ is called a reductive homogeneous space [10]. In fact, this is the case for the construction given in [1] for the MacDowell-Mansouri action. We shall extend this idea, and we shall see that Cartan geometries modeled on reductive homogeneous spaces are intrinsically linked to the notion of cosmological constant in specific cases. In the examples given in this paper, reductive homogeneous spaces will be implemented at the infinitesimal level through symmetric Lie algebras. Using the concept of mutation, two symmetric Lie algebras where $H$ is either the Lorentz or Lorentz×$\times$×Weyl group will be given as resulting from a mutation procedure on which a Cartan geometry will be constructed along the line of [11].
The paper is organized as follows. Section 2 will introduce and describe the gauge theories associated to the Lie groups $G=SO(4,1)$ (respectively $SO(3,2)$) for $\Lambda_{0}>0$ (respectively $\Lambda_{0}<0$) considered as "mutations" (according to [11]) of the group $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then, the gauge invariant action (8) will be studied for these Lie groups $G$ and $H$. Subsenquently, a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together. Results of Section 2 are adaptable to other spacetimes whether Riemannian or Galilean or for other Lie groups.
Section 3 will be devoted to the construction of a similar action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(3,1)=SO(3,1)\times(\mathbb{R}_{+}\setminus\{0\})$ the Lorentz×$\times$×Weyl group of signature (3,1)(3,1)( 3 , 1 ), so that the infinitesimal Klein geometry $(\mathfrak{g},\mathfrak{h})$ in the sense of [11] turns out to be a symmetric Lie algebra. Then, mutating the $(\mathfrak{g},\mathfrak{h})$ model allows to mimic what we did in the previous Lorentz case. Reducing the quotient $\mathfrak{g}/\mathfrak{h}$ to be of dimension 444 by constraining the "pair of frames" $(\alpha,\beta)$ one can retrieve an action comprising all terms described in the previous example + a kinetic term for a scalar field (dilations).
@@
Lastly, in Section 4, by studying invariant polynomials for reductive homogeneous spaces, it will be shown that (upon restricting to connections with compact supports on $\mathscr{M}$) the mutated geometries can be considered as deformations such that the gauge invariant actions constructed in the previous sections are asymptotically invariant under the choice of this type of connections, and thus lead to topological theories.
@@
2 Invariant polynomials for a Cartan connection with Lorentzian spacetimes
2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
@@
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive [11], that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant.
@@
$\eta=\begin{pmatrix}-1&0\\ 0&+I_{3}\end{pmatrix}\ .$
A group element $g\in G=ISO(3,1)$ is parametrized as ${\displaystyle g=\begin{pmatrix}\Lambda&a\\ 0&1\end{pmatrix}}$ where $a\in\mathbb{R}^{3,1}$. Linearization leads to:
@@
Definition 2.1.
Let $(\mathfrak{g},\mathfrak{h})$ and $(\mathfrak{g}^{\prime},\mathfrak{h})$ be two model geometries. A mutation map corresponds to an $Ad(H)$ module isomorphism $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ (i.e. $\mu(Ad_{h}(u))=Ad_{h}(\mu(u))\,\forall u\in\mathfrak{g}$) satisfying:
@@
$N=\begin{pmatrix}\eta&0\\ 0&-k_{1}/k_{2}\end{pmatrix}$
@@
$\left\{\begin{array}[]{lcll}\dfrac{k_{1}}{k_{2}}=-1\Leftrightarrow G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0,\\ &&&\\ \dfrac{k_{1}}{k_{2}}=+1\Leftrightarrow G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (5)
Throughout the paper $k_{1}$ and $k_{2}$ will be considered as real numbers. It is however of physical interest to consider the mutation parameters as scalar fields, thus dealing with point-wisely dependent mutations; this will be studied in a forthcoming paper. Other papers considering dynamical versions of physical constants in the context of gravity include for instance [12, 13].
On the other hand, the canonical decomposition of a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [14, Prop.2.1, Chapter XI] for more examples) corresponds to the commutators:
@@
[…]
@@
where we have used the commutators $[J_{ab},M_{c}]=\eta_{bc}M_{a}-\eta_{ac}M_{b}$ and $[M_{a},M_{b}]=k_{1}k_{2}\,J_{ab}$. [^7] It is interesting to notice that the product of the mutation parameters $k_{1}$ and $k_{2}$ can be related to a deformation parameter $\alpha$ as given in [15, eq.(2.1)].
@@
[…]
@@
2.2 Hodge *-operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, [^8] In this Section, one denotes $\mathfrak{g}$ the mutated algebra $\mathfrak{g}^{\prime}$ defined in (1) and $\mathfrak{p}\simeq\mathfrak{m}$. the Killing form $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})$ on $\mathfrak{g}$ splits into two parts:
[…]
where the trace is performed with respect to the generators $J_{ab},a<b$ and $M_{a}$. [^9] In full generality, one has $\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})=(m-1)\operatorname{Tr}(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{h}})+2k_{1}k_{2}(m-1)\eta(\varphi_{\mathfrak{p}},\varphi^{\prime}_{\mathfrak{p}})$.
@@ 2.2 Hodge *-operator
[…]
@@ 2.3 Associated action
The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+\dfrac{2k}{\ell^{2}}\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}=K_{\mathfrak{g}}(\varphi,\varphi^{\prime})/3$, see (2.2).
By combining linearly the invariant polynomials $\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ and $\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})$ (corresponding respectively to deformations of the Euler and Pontrjagin numbers of $R$ with deformation $k\xi$) with the Pontrjagin number $\mathdutchcal{P}$ (see (7)) associated to $\bar{\Omega}$, one can build the action (for $e,r,y$ real numbers [^10] Complex numbers may be taken for example to describe a complex formulation [16, p.13] of Loop Quantum Gravity (LQG) with a complex Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{e}{r+2y}$.):
@@
Using the isomorphism $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\rightarrow\mathfrak{spin}(n-1,1)$ between the Lorentz Lie algebra $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$, [17, p.192], whose action on the Lie algebra generators of $\mathfrak{so}(n-1,1)$ is given by:
@@
the covariant derivative associated to $\varpi_{\mathfrak{h}}$ on the space of spinors reads $D_{\mathfrak{h}}=d+\Phi_{\varepsilon}(\varpi_{\mathfrak{h}})=(\partial_{\mu}+\dfrac{1}{4}\varpi_{\mathfrak{h}\>\,ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$. One can then construct the following matter action: [^11] Since the metrics at the level of the Lie algebra and on $\mathscr{M}$ are always defined up to the factors $\kappa$ and $\zeta$, we set $\kappa^{2}\zeta=\tfrac{\ell^{4}}{216k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [17, p.197].
@@
where $\mathrm{Re}S_{D}$ corresponds to the real part of the Dirac action as described in [17, p.197] and $\tilde{*}$ is the Hodge star operator defined from $\tilde{g}$.
@@
While for general mutation $\mu$ with parameters $k,\ell$, (recall the particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (4) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
@@
$\qquad\quad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$
@@
For $y=0$ and $\alpha_{1}=-\dfrac{1}{16\pi G}$, one identifies [^13] At this stage a remark is in order. Indeed, the $[\mathfrak{m},\mathfrak{m}]$ bracket occurring in the $\mathfrak{h}$-part of the curvature, see (12) which is related to the mutation parameters $k,\ell$ turns out to be related to the bare cosmological constant $\Lambda_{0}$. This gives a physical interpretation of the mutation. See Section 4 where the topological features of the theory will be linked to the vanishing of the bare cosmological constant $\Lambda_{0}$. $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ and $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ such that applying the Hodge star * to the first equation (9) gives the Einstein equations with bare cosmological constant $\Lambda_{0}$ while (10) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:
@@
One may notice that the previous identification allows to recover the value (up to the mutation parameter $k$ and sign convention) $e/(32\pi^{2})=-\ell^{2}/(64\pi Gk)$ of the parameter $\alpha$ of [19, Sec II] in front of the Euler density built out of $R$, leading to finite Noether charges and a regularized Euclidean action as mentioned in [20, 19].
@@
  • If $\Lambda_{\text{vac}}<\Lambda_{exp}$, agreement with $\Lambda_{exp}$ would require a mutation parameter $k<0$. A particular case is $k=-1$ $\Leftrightarrow G=SO(4,1)$ (dS spacetime). Especially, $\rho_{\text{vac}}=0$ whenever one imposes normal ordering as mentioned in [23].
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this either implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkowski spacetime or $\Lambda_{0}=-\dfrac{3k}{\ell^{2}}\to 0$. In this limit, the contribution (12) to the $\mathfrak{h}$-part of the Cartan curvature becomes small and the theory can be considered as topological. This point will be discussed in detail in Section 4.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $y=0$. Even if they do not affect the classical equations of motion, the Euler and Pontrjagin densities of $R$ as well as the Nieh-Yan term (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [16, p.2]. The Nieh-Yan term also gives a non trivial contribution to the Barbero-Immirzi parameter $\gamma=\dfrac{e}{r+2y}$ via the parameter $r$.
@@
The main difference compared to the literature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the bare cosmological constant and by, respectively, the coupling constant of the Palatini action and the Barbero-Immirzi (B-I) parameter as well as $r$. Moreover, notice that one may identify the coefficient $\alpha_{3}=\dfrac{r+y}{8\pi^{2}}$ related to the Pontrjagin density built from $R$ with the parameter $\beta$ of [19, eq.(26)]. In this case one would then be able to identify (up to a sign convention) the last remaining physical parameter:

### v7

@@ Abstract
[…]
@@
I Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri (MM) gravity and BF theory, see e.g. [1]. They can be mathematically described in terms of Cartan geometries on a manifold $\mathscr{M}$ modeled on a homogeneous space $G/H$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their corresponding Lie algebras. When the geometry is reductive, namely, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$, a $\operatorname{Ad}(H)$-invariant decomposition, the associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$, one with values in $\mathfrak{h}$ and the other one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part is arranged to be $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda_{0}}{3\epsilon}}\begin{pmatrix}0&\beta\\ -\epsilon\bar{\beta}&0\end{pmatrix}$ and is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$) where $\bar{\beta}=\eta(\beta)$ with $\eta$ the Minkowski metric.
@@ I Introduction
Definition I.1.
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a $1$-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$, the Lie algebra of $G$, satisfying the conditions [2, p128]:
@@
[…]
@@
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$. This action is $\operatorname{Ad}(H)$-invariant, i.e. Lorentz invariant.
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$ originating from the $\mathfrak{so}(m)$-valued connection $1$-form $\omega_{\mathfrak{so}}$. This is an $SO(m)$-invariant polynomial.
For a $4$-dimensional manifold one has:
@@
where $\bar{\Omega}_{\mathfrak{h}}=R-\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$, and $\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}$ acts as the deformation of the $\bar{\Omega}_{\mathfrak{so}}=R$ curvature.
@@
[…]
The paper is organized as follows. Section II will introduce and describe the gauge theories associated to the Lie groups $G=SO(4,1)$ (respectively $SO(3,2)$) for $\Lambda_{0}>0$ (respectively $\Lambda_{0}<0$) considered as "mutations" (according to [11]) of the group $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then, the gauge invariant action (8) will be studied for these Lie groups $G$ and $H$. Subsenquently, a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together. Results of Section II are adaptable to other spacetimes whether Riemannian or Galilean or for other Lie groups.
[…]
@@
Lastly, in Section IV, by studying invariant polynomials for reductive homogeneous spaces, it will be shown that (upon restricting to connections with compact supports on $\mathscr{M}$) the mutated geometries can be considered as deformations such that the gauge invariant actions constructed in the previous sections are asymptotically invariant under the choice of this type of connections, and thus lead to topological theories.
@@
II Invariant polynomials for a Cartan connection with Lorentzian spacetimes
II.1 Lorentzian (dS and AdS) spacetimes as mutations of an
@@
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive [11], that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $\operatorname{Ad}(H)$-invariant.
@@
$\eta=\begin{pmatrix}-1&0\\ 0&+I_{3}\end{pmatrix}\$.
A group element $g\in G=ISO(3,1)$ is parametrized as $g=\begin{pmatrix}\Lambda&a\\ 0&1\end{pmatrix}$} where $a\in\mathbb{R}^{3,1}$. Linearization leads to:
@@
Definition II.1.
Let $(\mathfrak{g},\mathfrak{h})$ and $(\mathfrak{g}^{\prime},\mathfrak{h})$ be two model geometries. A mutation map corresponds to an $\operatorname{Ad}(H)$ module isomorphism $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ (i.e. $\mu(\operatorname{Ad}_{h}(u))=\operatorname{Ad}_{h}(\mu(u))\,\forall u\in\mathfrak{g}$) satisfying:
@@
$N=\begin{pmatrix}\eta&0\\ 0&-k_{1}/k_{2}\end{pmatrix}\$
@@
[…]
Throughout the paper $k_{1}$ and $k_{2}$ will be considered as real numbers. It is however of physical interest to consider the mutation parameters as scalar fields, thus dealing with point-wisely dependent mutations. This is studied in [12]. Other papers considering dynamical versions of physical constants in the context of gravity include for instance [13, 14].
On the other hand, the canonical decomposition of a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [15, Prop.2.1, Chapter XI] for more examples) corresponds to the commutators:
@@
$\tilde{e}^{b}(e_{a})=\delta_{~a}^{b}\text{ (dual basis)},\qquad\qquad\qquad$ $\bar{e}_{a}(e_{b})=\eta_{ab},$ $J_{ab}=e_{a}\bar{e}_{b}-e_{b}\bar{e}_{a}$
@@
where we have used the commutators $J_{ab},M_{c}]=\eta_{bc}M_{a}-\eta_{ac}M_{b}$ and $M_{a},M_{b}]=k_{1}k_{2}\,J_{ab}$. [^7] It is interesting to notice that the product of the mutation parameters $k_{1}$ and $k_{2}$ can be related to a deformation parameter $\alpha$ as given in [16, eq.(2.1)].
@@
$(D\bar{\Omega})_{\mathfrak{h}}$ $=dR+[A,R]+\dfrac{k}{\ell^{2}}\bigl(\underbrace{d(\beta\wedge\bar{\beta})+[A,\beta\wedge\bar{\beta}]+\beta\wedge\bar{T}-T\wedge\bar{\beta}}_{=\,0}\bigr)=dR+[A,R]=0$
@@
II.2 Hodge *-operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, [^8] In this Section, one denotes $\mathfrak{g}$ the mutated algebra $\mathfrak{g}^{\prime}$ defined in (9) and $\mathfrak{p}\simeq\mathfrak{m}$. the Killing form $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(\operatorname{ad}_{\varphi}\circ\operatorname{ad}_{\varphi^{\prime}})$ on $\mathfrak{g}$ splits into two parts:
$K_{\mathfrak{h}}:\mathfrak{h}\times\mathfrak{h}\rightarrow\mathbb{R}$ $\text{ and }K_{\mathfrak{m}}:\mathfrak{m}\times\mathfrak{m}\rightarrow\mathbb{R}$ (24)
where the trace is performed with respect to the generators $J_{ab},a<b$ and $M_{a}$. [^9] In full generality, one has $\operatorname{Tr}(\operatorname{ad}_{\varphi}\circ\operatorname{ad}_{\varphi^{\prime}})=(m-1)\operatorname{Tr}(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{h}})+2k_{1}k_{2}(m-1)\eta(\varphi_{\mathfrak{p}},\varphi^{\prime}_{\mathfrak{p}})$.
@@ II.2 Hodge *-operator
Let $\varepsilon_{a_{1}a_{2}...a_{m}}$ be the Levi-Civita symbol in dimension $m$, such that $\varepsilon_{12...m}=1$. From the metric $g$ and its determinant $|g|$, one can define a Hodge star operator * that acts on $\omega$ as:
@@ II.3 Associated action
The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+\dfrac{2k}{\ell^{2}}\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}=K_{\mathfrak{g}}(\varphi,\varphi^{\prime})/3$, see (II.2).
By combining linearly the invariant polynomials $\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ and $\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})$ (corresponding respectively to deformations of the Euler and Pontrjagin numbers of $R$ with deformation $k\xi$) with the Pontrjagin number $\mathdutchcal{P}$ (see (7)) associated to $\bar{\Omega}$, one can build the action (for $e,r,y$ real numbers [^10] Complex numbers may be taken for example to describe a complex formulation [17, p.13] of Loop Quantum Gravity (LQG) with a complex Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{e}{r+2y}$.):
@@
Using the isomorphism $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\rightarrow\mathfrak{spin}(n-1,1)$ between the Lorentz Lie algebra $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$, [18, p.192], whose action on the Lie algebra generators of $\mathfrak{so}(n-1,1)$ is given by:
@@
the covariant derivative associated to $\varpi_{\mathfrak{h}}$ on the space of spinors reads $D_{\mathfrak{h}}=d+\Phi_{\varepsilon}(\varpi_{\mathfrak{h}})=(\partial_{\mu}+\dfrac{1}{4}\varpi_{\mathfrak{h}\>\,ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$. One can then construct the following matter action: [^11] Since the metrics at the level of the Lie algebra and on $\mathscr{M}$ are always defined up to the factors $\kappa$ and $\zeta$, we set $\kappa^{2}\zeta=\tfrac{\ell^{4}}{216k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [18, p.197].
@@
where $\mathrm{Re}S_{D}$ corresponds to the real part of the Dirac action as described in [18, p.197] and $\tilde{*}$ is the Hodge star operator defined from $\tilde{g}$.
@@
While for general mutation $\mu$ with parameters $k,\ell$, (recall the particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (31) and $S_{G}$ (II.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
@@
$\qquad\quad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr)$
@@
For $y=0$ and $\alpha_{1}=-\dfrac{1}{16\pi G}$, one identifies [^13] At this stage a remark is in order. Indeed, the $\mathfrak{m},\mathfrak{m}]$ bracket occurring in the $\mathfrak{h}$-part of the curvature, see (20) which is related to the mutation parameters $k,\ell$ turns out to be related to the bare cosmological constant $\Lambda_{0}$. This gives a physical interpretation of the mutation. See Section IV where the topological features of the theory will be linked to the vanishing of the bare cosmological constant $\Lambda_{0}$. $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ and $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ such that applying the Hodge star * to the first equation (36) gives the Einstein equations with bare cosmological constant $\Lambda_{0}$ while (37) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:
@@
One may notice that the previous identification allows to recover the value (up to the mutation parameter $k$ and sign convention) $e/(32\pi^{2})=-\ell^{2}/(64\pi Gk)$ of the parameter $\alpha$ of [20, Sec II] in front of the Euler density built out of $R$, leading to finite Noether charges and a regularized Euclidean action as mentioned in [21, 20].
@@
  • If $\Lambda_{\text{vac}}<\Lambda_{exp}$, agreement with $\Lambda_{exp}$ would require a mutation parameter $k<0$. A particular case is k=1k=-1 $\Leftrightarrow G=SO(4,1)$ (dS spacetime). Especially, $\rho_{\text{vac}}=0$ whenever one imposes normal ordering as mentioned in [24].
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this either implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkowski spacetime or $\Lambda_{0}=-\dfrac{3k}{\ell^{2}}\to 0$. In this limit, the contribution (20) to the $\mathfrak{h}$-part of the Cartan curvature becomes small and the theory can be considered as topological. This point will be discussed in detail in Section IV.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $y=0$. Even if they do not affect the classical equations of motion, the Euler and Pontrjagin densities of $R$ as well as the Nieh-Yan term (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [17, p.2]. The Nieh-Yan term also gives a non trivial contribution to the Barbero-Immirzi parameter $\gamma=\dfrac{e}{r+2y}$ via the parameter $r$.
@@
The way (35) is built from invariant polynomials in the curvature with mutated geometries may be of use to obtain a generalization of the action described in the context of qeGR in [26] that would incorporate all the topological terms appearing in (35) and [17]. The main difference compared to the literature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the bare cosmological constant and by, respectively, the coupling constant of the Palatini action and the Barbero-Immirzi (B-I) parameter as well as $r$. This is due to how we constructed the gauge action from topological arguments. Moreover, notice that one may identify the coefficient $\alpha_{3}=\dfrac{r+y}{8\pi^{2}}$ related to the Pontrjagin density built from $R$ with the parameter $\beta$ of [20, eq.(26)]. In this case one would then be able to identify (up to a sign convention) the last remaining physical parameter: